Calculate the amount of heat required to raise the temperature of 225 g of water from 20.0 ⁰C to 100.0 ⁰C. _ Joules
1
Expert's answer
2019-05-14T03:40:12-0400

Solution.

"Q = c \times m \times (t2-t1)"


Q = 4200 * 0.225 * (100-20) = 75.6 kJ

Answer:

Q = 75.6 kJ